OpenStudy (anonymous):

Which statements are true? Choose all that are correct. A. –7b3(6b2 + 2) = –42b5 – 14b3 B. –2x2(4x3 + 5) = –8x6 – 10x2 C. –5y4(5y2 + 4) = –25y8 – 20y4 D. –3d3(3d2 + 2) = –9d5 – 6d3

OpenStudy (anonymous):

1)-7b3*6b2=-42b5 and -7b3*2=-14b3 so correct 2)-2x^2*4x^3=-8x^5 not -8x^6 3)-5y^4*5y^2=-25y6 4)-3d^3*3d^2=-9d5 and the last thing you can do it :)